OpenStudy (mathhelp346):

PLEASE HELP!! The perimeter of a triangle is 51 centimeters. Side b is 6 centimeters longer than side a. Side c is 5 centimeters shorter than 3 times the length of side a. Find the length of each side. Would the equation be something like: 51=b+6+3a+(5-c)

ganeshie8 (ganeshie8):

you need to form multiple equations, right ?

ganeshie8 (ganeshie8):

The perimeter of a triangle is 51 centimeters a+b+c = 51 -------------(1)

ganeshie8 (ganeshie8):

Side b is 6 centimeters longer than side a. b = a+6 -----------(2)

OpenStudy (mathhelp346):

b=6+a c=3a-5

ganeshie8 (ganeshie8):

Yes, Side c is 5 centimeters shorter than 3 times the length of side a. c=3a-5 --------------(3)

OpenStudy (mathhelp346):

Ok

ganeshie8 (ganeshie8):

3 equations, and 3 unknowns, we can solve them easily :)

ganeshie8 (ganeshie8):

simply substitute b & c values from equations (2) & (3) in equation (1)

OpenStudy (mathhelp346):

Ok I got a=10 b=16 c=25 Is this correct?

ganeshie8 (ganeshie8):

a+b+c = 51 a+ a+6 + 3a-5 = 51 5a+1 = 51 5a = 50 a = 10

ganeshie8 (ganeshie8):

Yes, they're right ! good work !!

OpenStudy (mathhelp346):

Thank you ^-^

ganeshie8 (ganeshie8):

np :)